OpenStudy (anonymous):

Which function below is the inverse of f(x) = The quantity of two x plus three, over five.?

OpenStudy (anonymous):

\[\huge f(x)=\frac{ 2x+3 }{ 5 }\]

OpenStudy (anonymous):

Step 1: Change f(x) to y. Step 2: Flip x and y. Step 3: Solve for y using the formula y=mx+b.

OpenStudy (anonymous):

Can you try it?

OpenStudy (anonymous):

\( \huge f(x)=\frac{ 2x+3 }{ 5 } \) Step 1: Change f(x) to y \( \huge y=\frac{ 2x+3 }{ 5 } \) Step 2: flip x and y \( \huge x=\frac{ 2y+3 }{ 5 } \) Step 3: Solve for y. Can you solve for y ??? \( \huge x=\frac{ 2y+3 }{ 5 } \)

OpenStudy (anonymous):

Solve for y \( \huge x=\frac{ 2y+3 }{ 5 } \) Lets get rid of the fraction. Times each side by 5 \( \huge 5 * x=\frac{ 2y+3 }{ 5 } * 5 \) \( \huge 5 x=2y+3 \) Can you solve for y in this equation \( \huge 5 x=2y+3 \) now ? \( \huge 5 x=2y+3 \)

OpenStudy (anonymous):

y=(5x-3)/2

OpenStudy (anonymous):

That is correct!!!

OpenStudy (anonymous):

So the inverse is \( \huge f(x)^{-1} = \frac{5x-3}{2}\)